OpenStudy (jojo12g):

What value of the variable makes the statement true? m + 4 = –2.8 A. –6.8 B. –1.2 C. 1.2 D. 6.8

OpenStudy (alexandervonhumboldt2):

m+4-4=-2.8-4 m=-(2.8+4)=6.8

OpenStudy (alexandervonhumboldt2):

woops -6.8

OpenStudy (anonymous):

-6.8+4=-2.8

OpenStudy (derpisthewordorisit):

A

OpenStudy (jojo12g):

thanks guys

OpenStudy (derpisthewordorisit):

It's A, solve for M.

OpenStudy (derpisthewordorisit):

It's really easy.

OpenStudy (derpisthewordorisit):

-2.8-4=-6.8, and for you to know to see if it's the right answer, just do -6.8+4=-2.8